# CoxSE: Exploring the Potential of Self-Explaining Neural Networks with Cox Proportional Hazards Model for Survival Analysis

### v1

@@ 6 Explanations Assessment on Real Datasets
In this section, we discuss the results of the two real datasets FLCHAIN and SEER. With the existence of co-linearities and correlations between the features, models can rely on different subsets of features which result in different explanations. However, as we do not know the ground truth of these datasets, an interesting validation method would be to compare how consistent the explanations provided by the model are with the explanations provided by an external method like SHAP applied to the model itself.
The SHAP method models the features’ contribution to the model’s decision. Whereas, CoxSE and CoxSENAM models produce explanations in terms of locally linear weights, which represent the sensitivity to change in the respective features. Based on such weights, we can compute the contribution in a similar way SHAP does for a linear model.
Given a linear model trained on a dataset with $m$ features and has weights $\beta_{i}:i\in 0..m-1$, the Shapley Values $\Phi_{i}$ of the feature $x_{i}$ weighted by $\beta_{i}$ can be written as:
@@
Based on Equation 27, we compute the contributions of the features for the CoxSE and the CoxSENAM by replacing the constant weights with their variable weights as:
$\Phi_{i}=w_{i}(\mathbf{x})(x_{i}-\mathbb{E}(x_{i}))$ (28)
We compared the rank order of the features provided by each model against the rank order of the SHAP explanation of the same model. We used the weighted-$\tau$ [31] which is a weighted version of Kendall-$\tau$ [32] rank correlation index. The weighted-$\tau$ penalizes the disagreement in the higher ranked items more than the lower ones. This makes it convenient for comparing feature ranking.
@@
Figure 11: Rank correlation between the models’ explanation and its own Shapley Values
Except for the approximation error of the SHAP method, the results in Figure 11 show almost a perfect correlation between the linear CPH model’s explanations and its Shapley Values. This is understandable due to the direct relationship between Shapley Values and the linear weights as shown in Equation 27. The results also show a clear advantage of the SENN-based models (CoxSE and CoxSENAM) which have much higher correlations with their own Shapley Values than the CoxNAM model. This can be attributed to the local linearity of the models encouraged by the regularization term $\mathcal{L}_{2}$.
It is worth noting that there was a complete match between the ranking of the aggregated explanations of all models with the ranking of their own aggregated Shapley Values except for the case of CoxNAM on the Lin dataset shown in Figure 4(d). This can explain the lower correlation between CoxNAM and its SHAP explanations in the case of the real datasets. Moreover, this strengthens the hypothesis that transforming the function into an additive function does not guarantee valid explanations and highlights the importance of the regularization terms in the SENN-based models in the validity of the generated explanations.
@@ 7 Conclusion
In this work, we explored how self-explaining neural networks with the CPH model can be adapted for survival modeling. We proposed two CPH-based models, CoxSE relying on the SENN structure, and CoxSENAM which is a hybrid model between SENN and the NAM structures. The two models share the same type of output and loss function providing local explanations as a locally linear approximation of the underlying function. In this regard, unlike Shapley Values and the CoxNAM model which provide explanations as feature contributions, our models’ explanations come in the form of local weights. As such, they highlight the sensitivity of the outcome to changes in the features in the vicinity of the point of interest. However, we illustrated that feature contributions can be computed directly from such weights similar to the case of a linear model, however, with variable weights.
@@
We performed several experiments with synthetic and real datasets which highlighted the properties of the proposed models. In terms of performance, CoxSE was shown to be a flexible model that matches the performance of its black-box counterpart; DeepSurv. However, similar to the CoxNAM model, CoxSENAM is less flexible due to the NAM structure which cannot handle interactions between the features. Nevertheless, CoxSE and CoxSENAM showed better stability than CoxNAM and DeepSurv (explained by SHAP) and better matching with the ground truth than CoxNAM. This is due to the regularization term weighted by $\alpha$ in their loss function that directly encourages local similarity of explanations.
@@
On the other hand, the NAM-based models showed better robustness to noisy features than the CoxSE model. This can be attributed to the structure of the NAM models where the noise contribution to the output comes as additive noise. This makes it easier for the NAM-based models to isolate the noise transforming it into a constant value added to the output. However, our hybrid model, CoxSENAM, showed better robustness than CoxNAM, assigning less importance to noisy features. This can be attributed to the second regularization term weighted by $\beta$ encouraging the noise contribution to be close to zero.
@@
Lastly, using real datasets, our models CoxSE and CoxSENAM demonstrated better consistency between their intrinsic explanations and the post-hoc explanation provided by SHAP. This was also observed on the synthetic datasets which arguably indicates the validity of such explanations in describing the underlying contributions of the features. On the other hand, CoxNAM demonstrated less consistency with its post-hoc explanations on both, the real and synthetic datasets.
@@ Appendix A Hyperparameters Tuning
Five-fold cross-validation was used to tune the models’ hyperparameters number of layers, number of nodes per layer, $L_{2}$ regularization weight, dropout rate, and learning rate. Moreover, we tuned $\alpha$ and $\beta$ weights of the loss function for the CoxSE and CoxSENAM models. The selected hyperparameters are listed in Tables 2345, and 6.
Hyperparameter CoxSE CoxSENAM DeepSurv CoxNAM CPH
No. Layers 4 1 3 2 -
No. Nodes 128 4 4 32 -
L2 0 0.5 0.1 0.7 -
Dropout 0.4 0 0 0 -
Learning Rate 0.01 0.1 0.05 0.1 0.1
alpha 0.5 0.5 - - -
beta 0 0.005 - - -
Table 2: hyperparameters of the models on the Lin dataset
Hyperparameter CoxSE CoxSENAM DeepSurv CoxNAM CPH
No. Layers 1 1 1 4 -
No. Nodes 4 128 128 128 -
L2 0.5 1 0.01 1 -
Dropout 0 0.4 0 0 -
Learning Rate 0.001 0.006 0.001 0.0009 0.0003
alpha 3e-6 9e-6 - - -
beta 5e-4 9e-6 - - -
Table 3: hyperparameters of the models on the NonLin dataset
Hyperparameter CoxSE CoxSENAM DeepSurv CoxNAM CPH
No. Layers 1 1 1 2 -
No. Nodes 32 64 128 128 -
L2 0.5 0 0.5 0.05 -
Dropout 0 0.1 0 0 -
Learning Rate 0.1 0.1 0.0001 0.05 0.03
alpha 1e-5 3e-3 - - -
beta 1e-4 5e-6 - - -
Table 4: hyperparameters of the models on the NonLinX dataset
Hyperparameter CoxSE CoxSENAM DeepSurv CoxNAM CPH
No. Layers 4 4 1 4 -
No. Nodes 4 32 16 64 -
L2 0.5 0.001 1 0 -
Dropout 0.3 0.4 0.1 0 -
Learning Rate 0.01 0.003 0.0005 0.001 0.09
alpha 4e-5 6e-5 - - -
beta 6e-3 5e-5 - - -
Table 5: hyperparameters of the models on the FLCHAIN dataset
Hyperparameter CoxSE CoxSENAM DeepSurv CoxNAM CPH
No. Layers 1 4 1 4 -
No. Nodes 4 16 32 16 -
L2 0.5 0.5 1 0.05 -
Dropout 0 0 0.2 0 -
Learning Rate 0.0009 0.001 0.0005 0.0005 0.01
alpha 1e-5 7e-5 - - -
beta 3e-5 5e-6 - - -
Table 6: hyperparameters of the models on the SEER dataset

### v2

@@ 6 Explanations Assessment on Real Datasets
In this section, we discuss the results of the two real datasets, FLCHAIN and SEER. With the existence of collinearities and correlations between the features, models can rely on different subsets of features, which result in different explanations. However, as we do not know the ground truth of these datasets, an interesting validation method would be to compare how consistent the explanations provided by the model are with the explanations provided by an external method like SHAP applied to the model itself.
The SHAP method computes the features’ contribution to the model’s decision. Whereas, CoxSE and CoxSENAM models produce explanations in terms of locally linear weights, which represent the sensitivity to change in the respective features. Based on such weights, we can compute the contribution in a similar way to how SHAP does for a linear model.
Given a linear model trained on a dataset with $m$ features and has weights $\beta_{i}:i\in 0..m-1$, the Shapley Value $\Phi_{i}$ of the feature $x_{i}$ weighted by $\beta_{i}$ can be written as:
@@
Based on Eq. 27, we compute the contributions of the features for the CoxSE and the CoxSENAM by replacing the constant weights with their variable weights as:
$\Phi_{i}=w_{i}(\bm{x})(x_{i}-\mathbb{E}(x_{i}))$ (28)
We compared the rank order of the aggregated feature importance provided by each model against the rank order of the aggregated SHAP explanation of the same model. We used the weighted-$\tau$ [Vigna2015], which is a weighted version of $\tau [Kendall$1938] rank correlation index. The weighted-$\tau$ penalizes the disagreement in the higher-ranked items more than the lower ones. This makes it convenient for comparing feature rankings.
@@
Figure 13: Rank correlation between the models’ explanations and their own Shapley Values
Except for the approximation error of the SHAP method, the results in Figure 13 show almost a perfect correlation between the linear CPH model’s explanations and its Shapley Values. This is understandable due to the direct relationship between Shapley Values and the linear weights as shown in Eq. 27. The results also show a clear advantage of the SENN-based models (CoxSE and CoxSENAM), which have much higher correlations with their own Shapley Values than the CoxNAM model. This can be attributed to the local linearity of the models encouraged by the regularization term $\mathcal{L}_{2}$.
It is worth noting that, on the synthetic datasets, there was a perfect match between the ranking of the aggregated explanations of all models with the ranking of their own aggregated Shapley Values, except for the case of CoxNAM on the Lin dataset shown in Figure 4(d). This can explain the lower correlation between CoxNAM and its SHAP explanations in the case of the real datasets. Moreover, this strengthens the hypothesis that transforming the function into an additive function does not guarantee valid explanations and highlights the importance of the regularization terms in the SENN-based models in the validity of the generated explanations.
6.1 Case Study: SEER Breast Cancer Dataset
To demonstrate the practical utility of our proposed models, we present a case study using the SEER breast cancer dataset. This dataset contains comprehensive clinical features from a large cohort of breast cancer patients, including tumor size, lymph node involvement, extension, and other pathological variables. We applied both CoxSE and CoxNAMSE models to analyze feature relevance across the patient population. Both proposed models consistently identified the same top five prognostic factors: tumor size, number of lymph nodes containing tumor, number of lymph nodes removed and examined, tumor extension, and the highest specific lymph node chain involved, as illustrated in Figure 14.
(a) CoxSE
(b) CoxSENAM
Figure 14: The top 10 important features of the SEER dataset identified by CoxSE and CoxSENAM models
These variables correspond closely to the core anatomical and pathological components of the TNM classification system, where tumor size and extension define the T category and lymph node involvement defines the N category [giuliano_eighth_2018], reflect the prognostic emphasis found in Nottingham Prognostic Index (NPI) [galea_nottingham_1992], and are consistent with insights reported in the literature [el_saghir_effects_2006]. This consistency lends support to the clinical plausibility of the model-generated explanations and suggests their potential utility in supporting clinical understanding and interpretation of survival outcomes, particularly in settings where transparency and interpretability are important. When compared with the baseline approaches (Figure 15), including CoxNAM, DeepSurv + SHAP, CPH, and RSF, the proposed models showed a noticeably different pattern in how they explained predictions. While the baseline models often ranked some demographic factors (such as age and race) and tumor grade among the top five features, the proposed models focused mainly on TNM-related variables. Moreover, the proposed methods exhibited greater differentiation in feature importance magnitudes, showing a clearer separation between dominant and secondary predictors than observed in other methods.
(a) CoxNAM
(b) DeepSurv + SHAP
(c) CPH
(d) RSF
Figure 15: The top 10 important features of the SEER dataset identified by CoxNAM, DeepSurv, CPH, and RSF models
@@ 7 Conclusion
In this work, we explored how self-explaining neural networks with the CPH model can be adapted for survival modeling. We proposed two CPH-based models, CoxSE relying on the SENN structure, and CoxSENAM, which is a hybrid model between the SENN and the NAM structures. The two models share the same type of output and loss function, providing local explanations as a locally linear approximation of the underlying function. In this regard, unlike Shapley Values and the CoxNAM model, which provide explanations as feature contributions, our models’ explanations come in the form of local weights. As such, they highlight the sensitivity of the outcome to changes in the features in the vicinity of the point of interest. However, we illustrated that feature contributions can be computed directly from such weights, similar to the case of a linear model, however, with variable weights.
@@
We performed several experiments with synthetic and real datasets, which highlighted the properties of the proposed models. In terms of performance, CoxSE was shown to be a flexible model that matches the performance of its black-box counterpart, DeepSurv. However, similar to the CoxNAM model, CoxSENAM is less flexible due to the NAM structure, which cannot handle interactions between the features. Nevertheless, CoxSE and CoxSENAM showed better stability than CoxNAM and DeepSurv (explained by SHAP) and better matching with the ground truth than CoxNAM. This is due to the regularization term weighted by $\alpha$ in their loss function that directly encourages local similarity of explanations.
@@
On the other hand, the NAM-based models showed better robustness to noisy features than the CoxSE model. This can be attributed to the structure of the NAM models, where the noise contribution to the output comes as additive noise. This makes it easier for the NAM-based models to isolate the noise, transforming it into a constant value added to the output. However, our hybrid model, CoxSENAM, showed better robustness than CoxNAM, assigning less importance to noisy features. This can be attributed to the second regularization term, weighted by $\beta$, encouraging the noise contribution to be close to zero.
@@
Lastly, using real datasets, our models, CoxSE and CoxSENAM, demonstrated better consistency between their intrinsic explanations and the post-hoc explanation provided by SHAP. This was also observed on the synthetic datasets, which arguably indicates the validity of such explanations in describing the underlying contributions of the features. On the other hand, CoxNAM demonstrated less consistency with its post-hoc explanations on both the real and synthetic datasets.
This work investigated how to employ the SENN model in Survival Analysis, relying on the Cox Proportional Hazards assumptions and using the partial log-likelihood objective function. This poses a restriction on the application of such a model to non-proportional hazards type of data. In non-proportional hazards scenarios, relying on models that assume proportionality can result in biased risk estimates and misleading interpretations of covariate effects over time. Consequently, the use of such models in these contexts may compromise both the validity of the predictions and the reliability of the derived explanations, as highlighted in [Schemper92, Alabdallah2022]. However, other survival model formulations can be used, such as direct regression of survival time or estimating the Probability Mass Function (PMF) in a discrete-time manner, which we leave for future work. Moreover, in this work, we assumed the original feature space to be explainable, which is the most common form of survival data. Consequently, we discarded the interpretable feature extraction part of the network (the autoencoder) along with its reconstruction loss function. However, in some cases, where the input space is not interpretable, like the case of image input, the feature-extraction part can be brought back, which is out of the scope of this work.
@@ Appendix A Hyperparameters Tuning
We performed hyperparameter tuning using five-fold cross-validation. The final hyperparameter configurations for all models are presented in Tables 3 9.
Table 3: hyperparameters of the CoxSE model
Hyperparameter Lin NonLin NonLinX Flchain SEER
No. Layers 4 1 1 4 1
No. Nodes 128 4 32 4 4
L2 0 0.5 0.5 0.5 0.5
Dropout 0.4 0 0 0.3 0
Learning Rate 0.01 0.001 0.1 0.01 0.0009
alpha 0.5 3e-6 1e-5 4e-5 1e-5
beta 0 5e-4 1e-4 6e-3 3e-5
Table 4: hyperparameters of the CoxSENAM model
Hyperparameter Lin NonLin NonLinX Flchain SEER
No. Layers 1 1 1 4 4
No. Nodes 4 128 64 32 16
L2 0.5 1 0 0.001 0.5
Dropout 0 0.4 0.1 0.4 0
Learning Rate 0.1 0.006 0.1 0.003 0.001
alpha 0.5 9e-6 3e-3 6e-5 7e-5
beta 0.005 9e-6 5e-6 5e-5 5e-6
Table 5: hyperparameters of the CoxNAM model
Hyperparameter Lin NonLin NonLinX Flchain SEER
No. Layers 2 4 2 4 4
No. Nodes 32 128 128 64 16
L2 0.7 1 0.05 0 0.05
Dropout 0 0 0 0 0
Learning Rate 0.1 0.0009 0.05 0.001 0.0005
Table 6: hyperparameters of the DeepSurv model
Hyperparameter Lin NonLin NonLinX Flchain SEER
No. Layers 3 1 1 1 1
No. Nodes 4 128 128 16 32
L2 0.1 0.01 0.5 1 1
Dropout 0 0 0 0.1 0.2
Learning Rate 0.05 0.001 0.0001 0.0005 0.0005
Table 7: hyperparameters of the CPH model
Hyperparameter Lin NonLin NonLinX Flchain SEER
Learning Rate 0.1 0.0003 0.03 0.09 0.01
Table 8: hyperparameters of the RSF model
Hyperparameter Lin NonLin NonLinX Flchain SEER
min_samples_split 20 10 10 10 10
min_samples_leaf 5 15 15 15 20
max_features log2 sqrt sqrt sqrt log2
max_depth 10 10 10 10 20
Table 9: hyperparameters of the DeepHit model
Hyperparameter Lin NonLin NonLinX Flchain SEER
No. bins 500 200 1000 1000 200
alpha 1e-1 1e-2 1e-2 5e-2 1e-1
sigma 9e-1 1.0 1e-3 9e-1 5e-1
Dropout 0.4 0 0.1 0.2 0.4
No. Layers 2 2 2 2 2
No. Nodes 32 128 16 64 64
Learning Rate 0.03 0.004 0.0009 0.0005 0.003